# Long and Short-Term Constraints Driven Safe Reinforcement Learning for Autonomous Driving

### v1

@@ V-A Experiment Settings
The MetaDrive [45] simulator is used as the training and evaluation environment in this paper since it has realistic vehicle dynamics and can provide various road maps which are randomly composed from elementary components such as the straight road, curve, intersection, roundabout, ramp, and fork. Each road map has a start point and a destination, connected with a route in the form of a series of navigation points. In our experiments, the AD vehicle’s task is to navigate from the start point to the destination based on navigation points.
@@ V-C Safety performance analysis compared with SOTA methods
TABLE III shows that our method has the lowest episode cost, which is much lower than TD3-Lag. Compared with TRPO-Lag and FAC, the best methods in classic CMDP and SafeRL-kit, our LSTC method achieves lower episode cost values by $0.26$ and $1.98$, respectively. The main reason is that the short-term constraint in the LSTC method considers the vehicle’s state-based safety, which encompasses avoidance of collisions with other vehicles, cones, obstacles, and warning signs, etc., so as to decrease the episode cost and improve the safety performance. In contrast, other safe RL methods rely on the goal-oriented constraint that only considers the long-term safety, achieving higher episode cost values. Since PPO doesn’t consider risks and aims to solely maximize the expected reward, its episode cost is the highest in all the comparative methods. The episode cost of REC is also much higher than our method because it learns the risk policy to minimize the cost. However, the risk policy takes over the main policy in hazardous states, which it struggles to learn.
@@ V-D Comparative results in different complex scenarios
We select the best method from each of the three kinds of methods including the unconstrained RL method, classic CMDP-based RL method, and safe RL method in SafeRL-kit, which are the PPO, PPO-Lag, and EPO, respectively, as the representative comparative methods. Each selected method is evaluated 50 times in the four scenarios. These ego vehicles randomly choose start and exit points on the roads with three lanes. Other traffic vehicles and static obstacles such as cones, roadblocks, and warning signs are randomly placed on the road. The success rate and average cost results are shown in TABLE IV. Additionally, we record the vehicle’s trajectories and the coordinates where accidents occur in all the episodes, which are as shown in Fig. 4.
TABLE IV shows that our method exhibits the highest success rates in all the scenarios and the lowest episode costs in the inputting ramp and sharp left-turn scenarios. According to the trajectory points shown in Fig. 4, The trajectory points of LSTC are evenly distributed across all scenarios, indicating that the vehicle maintains a consistent speed and mainly navigates in the middle lane with minimal lane changes. In the roundabout, collisions are relatively more frequent than other scenarios. The collisions primarily concentrate at the entrances and exits, where the trajectories are curved, because other vehicles easily collide with the ego vehicle while passing through the curves. The trajectory points of PPO are the sparsest of all, which indicates that the ego vehicle consistently maintains a high-speed navigation and disregards potential dangers. EPO uses the state-wise constraints for the cumulative cost as same as FAC to ensure strict safety guarantee, but EPO reduces the state-dependent Lagrange multipliers in FAC to a single fixed hyper-parameter. In this case, EPO achieves an overall lower episode cost, but its success rate is lower than other methods because the vehicle often chooses to depart from the lane to avoid collisions with other vehicles.

### v2

@@ V-A Experiment Settings
The MetaDrive [46] simulator is used as the training and evaluation environment in this paper since it has realistic vehicle dynamics and can provide various road maps which are randomly composed from elementary components such as the straight road, curve, intersection, roundabout, ramp, and fork. Each road map has a start point and a destination, connected with a route in the form of a series of navigation points. It is the most popular benchmark in recent studies about safe RL in autonomous driving such as [47, 48]. In our experiments, the AD vehicle’s task is to navigate from the start point to the destination based on navigation points.
@@ V-C Safety performance analysis compared with SOTA methods
TABLE III shows that our method has the lowest episode cost in all comparative methods, which means the fewest collisions and the safest decision in our method since the cost value is calculated by the numbers of collisions with other objects and departing from the lane. Compared with TRPO-Lag and FAC, the best methods in classic CMDP and SafeRL-kit, our LSTC method achieves lower episode cost values by $0.26$ and $1.98$, respectively. The main reason is that the short-term constraint in the LSTC method considers the vehicle’s state-based safety, which encompasses avoidance of collisions, so as to decrease the episode cost and improve the safety performance. In contrast, other safe RL methods rely on the goal-oriented constraint that only considers the long-term safety, achieving higher episode cost values. Since PPO doesn’t consider risks and aims to solely maximize the expected reward, its episode cost is the highest in all the comparative methods. The episode cost of REC is also much higher than our method because it learns the risk policy to minimize the cost. However, the risk policy takes over the main policy in hazardous states, which it struggles to learn.
@@ V-D Comparative results in different complex scenarios
We select the best method from each of the three kinds of methods including the unconstrained RL method, classic CMDP-based RL method, and safe RL method in SafeRL-kit, which are the PPO, PPO-Lag, and EPO, respectively, as the representative comparative methods. Each selected method is evaluated 50 times in the four scenarios. The ego vehicles randomly choose start and exit points on the roads with three lanes. Other traffic vehicles and static obstacles such as cones, roadblocks, and warning signs are randomly placed on the road. The success rates and average costs are shown in TABLE IV. Additionally, we record the vehicle’s trajectories and the coordinates where accidents occur in all the episodes, which are as shown in Fig. 4.
@@
TABLE IV shows that our method exhibits the highest success rates in all the scenarios and the lowest episode costs in the inputting ramp and sharp left-turn scenarios. According to the trajectory points shown in Fig. 4, The trajectory points of LSTC are evenly distributed across all scenarios, indicating that the vehicle maintains a consistent speed and mainly navigates in the middle lane with minimal lane changes. The trajectory points of PPO are the sparsest of all, which indicates that the ego vehicle consistently maintains a high-speed navigation and disregards potential dangers. EPO uses the state-wise constraints for the cumulative cost as same as FAC to ensure strict safety guarantee, but EPO reduces the state-dependent Lagrange multipliers in FAC to a single fixed hyper-parameter. In this case, EPO achieves an overall lower episode cost, but its success rate is lower than other methods because the vehicle often chooses to depart from the lane to avoid collisions with other vehicles. In the roundabout and intersection scenarios, the episode costs of our method are a slightly higher than those of EPO. The increased costs come from the collisions that mainly occur at the entrances and centers of roundabouts and intersections because the decisions made by LSTC are more cautious compared with EPO, leading to collides with other non-intelligent vehicles for the ego vehicle in LSTC when waiting to yield.